A portfolio has a market beta of 1.5, a factor loading of -2 to SMB and a sensitivity of 1 to HML. The portfolio has earned 22% return. What was the abnormal return on the portfolio, if the risk free rate was 0%, market's excess return was 6%, the SMB factor premium was -2% and the HML factor premium was 3%? 16% 6% 15% 21.5%
Answer #1

Abnormal return = Earned Return - Required return.

Firstly, calculate the required return,

Required return = (1.5 * 0.05) + (-2* -0.02) + (1 * 0.03)

Required return = 0.09 +0.04 + 0.03

Required return = 0.16

Thus,

Abnormal Return = Earned return - Required return

Abnormal Return = 0.22- 0.16

Abnormal Return = 0.06

Abnormal Return = 6%

Hope this will help, please do comment if you need any further explanation. Your feedback would be highly appreciated.